Identify the solution(s) of the system of equations, if any.
3x-2y=6
5x=5-25y

Answer by mananth(16946)

3.00 x -2.00 y = 6.00 .............1
Total value
5.00 x + 25.00 y = 5.00 .............2
Eliminate y
multiply (1)by 25.00
Multiply (2) by 2.00
75.00 x -50.00 y = 150.00
10.00 x + 50.00 y = 10.00
Add the two equations
85.00 x = 160.00
/ 85.00
x = 1.88
plug value of x in (1)
3.00 x -2.00 y = 6.00
5.65 -2.00 y = 6.00
-2.00 y = 6.00 -5.65
-2.00 y = 0.35
y = -0.18